OpenStudy (anonymous):

What is the sum of a 6-term geometric series if the first term is 8 and the last term is 134,456 ? Answers to choose from. a)129,513 b)130,624 c)156,864 d)172,846

OpenStudy (cwrw238):

let the first term = a then 6th term =ar^5 where r = common ration can you find r from this information?

OpenStudy (anonymous):

i might need some guidance but let me try!

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

i really dont know how i got this but i ended up with 832 as the common ratio.

OpenStudy (cwrw238):

6tt term = 134456 = ar^5 ist term = 8 so ar^5/a = r^5 = 134456/8 = 16807

OpenStudy (cwrw238):